OpenStudy (anonymous):

Find the Maclaurin series for ln(1+2x). For what values of x does the series converge? pls solution :D

OpenStudy (tkhunny):

Have you met this method? (You may wish to reference "Uniform Convergence") \(\dfrac{d}{dx}ln(1+2x) = \dfrac{2}{1+2x} = 2(1 - 2x + (2x)^2 - (2x)^{3} + ...)\) Then, \(ln(1+2x) = \dfrac{2}{1}x - \dfrac{4}{2}x^{2} + \dfrac{8}{3}x^3 - \dfrac{16}{4}x^{4} + \dfrac{32}{5}x^{5} - ... +(-1)^{n-1}\dfrac{2^{n}}{n}x^{n} + ...\) Sometimes, either the derivative or the antiderivative is easier! You have to do the convergence part! Ratio test?

OpenStudy (anonymous):

thanks big help :D